OpenStudy (karatechopper):

Find the measures of an exterior angle and an interior angles given the number of sides of each regular polygon. Round to the nearest tenth, if necessary. Number of sides: 16

OpenStudy (karatechopper):

So I started out by doing (n-2)(180) And I substituted n for 16 because n is the number of sides. so.. (16-2)(180) 14(180) =2520

OpenStudy (karatechopper):

And then I tried to find out what the measure of the interior angles was.

OpenStudy (ksaimouli):

Int angle=(n-2)* 180/n

OpenStudy (ksaimouli):

ext angle=360/n

OpenStudy (anonymous):

Now just divide by the number of sides

OpenStudy (karatechopper):

Yes that!! @ksaimouli I thought about that!!

OpenStudy (karatechopper):

R u sure? @zaynahf for some reason I thought it was wrong to do that.

OpenStudy (ksaimouli):

then u forgot to divide /n or number of sides in this case 16 i think

OpenStudy (karatechopper):

Lemme seee

OpenStudy (karatechopper):

2520/16=157.5 So are you saying that 157.5 is the measure of an interior angle?

OpenStudy (anonymous):

*SOrry, not the # of sides

OpenStudy (anonymous):

16-2= 14 2520/14

OpenStudy (karatechopper):

And then.. 360/16 =22.5 for exterior angle?

OpenStudy (karatechopper):

Oh! 14 for the triangles?

OpenStudy (karatechopper):

@zaynahf

OpenStudy (karatechopper):

That makes each side become 180 for interior!?! whaaa...makes no sense..

OpenStudy (anonymous):

I know, lol i just messed up

OpenStudy (karatechopper):

oh..so then how do I do this?

OpenStudy (anonymous):

Sorry, its been a long time -_- Here, i found this: Exterior angles always add up to 360 no matter how many sides. So to find an interior angle, it's easiest to find the exterior angle and subtract it from 180 since interior + exterior = 180 So 1 16-gon: 360 / 16 = 22.5; 180 - 22.5 = 157.5 degrees

OpenStudy (anonymous):

You were right :)

OpenStudy (karatechopper):

yes!!!

OpenStudy (anonymous):

Ok good :P Sorry again.. i forgot that stuff

OpenStudy (karatechopper):

Lol its ok!